# HyperGaussians: High-Dimensional Gaussian Splatting for High-Fidelity Animatable Face Avatars

### v2

@@ 3.2 HyperGaussians
This section introduces HyperGaussians, our proposed replacement for vanilla 3D Gaussians. HyperGaussians are an extension of 3D Gaussians to higher dimensions. As an example, a vanilla Gaussian in Kerbl et al. [18] has $m=3$ dimensions for the mean $\boldsymbol{\mu}$ representing its position. Intuitively, our HyperGaussian generalizes the vanilla Gaussian primitive to $(m+n)$ dimensions. We call $m$ the attribute dimensionality and the additional dimensions $n$ the latent dimensionality.
@@
This paragraph describes the HyperGaussian more formally. Consider a random vector $\boldsymbol{\gamma}\sim\mathcal{N}\bigl{(}\boldsymbol{\mu},\boldsymbol{\Sigma}\bigr{)}$ that is partitioned into vectors $\boldsymbol{\gamma}=(\boldsymbol{\gamma}_{a},\boldsymbol{\gamma}_{b})^{\top}$, where $\boldsymbol{\gamma}_{a}\in\mathbb{R}^{m}$ and $\boldsymbol{\gamma}_{b}\in\mathbb{R}^{n}$.
@@
with $\boldsymbol{\Sigma}_{ba}=\boldsymbol{\Sigma}_{ab}^{\top}$. In practice, $\boldsymbol{\gamma}_{a}$ is instantiated with 3D Gaussian attributes (e.g., position) and $\boldsymbol{\gamma}_{b}$ with a latent code. Following NDGS [10], we parameterize the mean $\boldsymbol{\mu}$ directly and decompose the covariance $\boldsymbol{\Sigma}$ into its Cholesky factor $\boldsymbol{L}$, such that $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}$, where $\boldsymbol{L}$ is a lower triangular matrix with positive diagonal entries. Training requires optimizing the parameters of the Cholesky factor $\boldsymbol{L}$ together with the mean $\boldsymbol{\mu}$. Please read on to the end of this section for a detailed explanation of which parameters require training.
@@
HyperGaussians can be splatted with the differentiable rasterizer proposed by Kerbl et al. [18] after being reduced to the attribute dimensionality. This can be done by conditioning on the latent dimensions. Geometrically, conditioning corresponds to taking an $m$-dimensional slice through the multivariate Gaussian with $(m+n)$ total dimensions. Fig. 4 shows two examples for $m=n=1$. More formally, we are interested in the conditional distribution $p(\boldsymbol{\gamma}_{a}|\boldsymbol{\gamma}_{b})=\mathcal{N}\bigl{(}\boldsymbol{\mu}_{a|b},\boldsymbol{\Sigma}_{a|b}\bigr{)}$. Since our HyperGaussian follows a multivariate Gaussian distribution, the conditional distribution can be computed in closed form [1]:
@@
The output of this computation is then used to recover a vanilla 3D Gaussian and can be splatted with the differentiable rasterizer introduced by Kerbl et al. [18].
Figure 4: Gaussian Conditioning on two examples with large (left) and small (right) uncertainty at different realizations of $\boldsymbol{\gamma}_{b}$. As a result, the conditional mean shifts, while the conditional covariance is the same for both slices.
@@
A naïve implementation of the conditioning in Eq. 4 is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$. The bottleneck lies in storing and inverting the conditional covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$. The latent dimensionality ($n$) is typically much larger than the dimensionality after conditioning $m$ ($n\gg m$) and therefore expensive to invert.
@@
To save the cost of this inversion, we apply the inverse covariance trick. The key idea is to reformulate the HyperGaussians in terms of their precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ such that $\boldsymbol{\gamma}\sim\mathcal{N}\bigl{(}\boldsymbol{\mu},\boldsymbol{\Lambda}^{-1}\bigr{)}$. Let’s consider the general block matrix view (as in Eq. 3)
with $\boldsymbol{\Lambda}_{ba}=\boldsymbol{\Lambda}_{ab}^{\top}$. The conditional mean and covariance matrix can now be expressed as
$\begin{split}\boldsymbol{\mu}_{a|b}&=\boldsymbol{\mu}_{a}-\boldsymbol{\Lambda}_{aa}^{-1}\boldsymbol{\Lambda}_{ab}\bigl{(}\boldsymbol{\gamma}_{b}-\boldsymbol{\mu}_{b}\bigr{)}\\ \boldsymbol{\Sigma}_{a|b}&=\boldsymbol{\Lambda}_{aa}^{-1}.\end{split}$ (6)
This new formulation only requires storing and inverting the much smaller block $\boldsymbol{\Lambda}_{aa}$ of the precision matrix. While this change seems minor in the derivation, it heavily improves both the speed and memory during runtime.
Optimizable Parameters
Eq. 6 is also helpful for understanding which parameters are optimized during training. It is not necessary to optimize the full precision matrix $\boldsymbol{\Lambda}$. The HyperGaussian parameters that require optimization are the means $\boldsymbol{\mu}_{a}$, $\boldsymbol{\mu}_{b}$ and the factors $\boldsymbol{L}_{11}$ and $\boldsymbol{L}_{21}$ of $\boldsymbol{\Lambda}_{aa}$ and $\boldsymbol{\Lambda}_{ab}$ coming from the Cholesky decomposition of $\boldsymbol{\Lambda}$. In short, the inverse covariance trick not only saves memory and computation but also reduces the number of optimizable parameters. Please see Sec. 4.3 for a benchmark comparison and Sec. A.1 for more details.
@@
Figure 5: Benchmark Results on conditioning for 14’876 HyperGaussians with conditional attribute dimension $m=3$ (e.g., position) and varying latent dimension $n$. We average the measurements across 1000 runs after initial warm-up. The point measurements were conducted on an NVIDIA GeForce RTX 2080 Ti 11GB. The benchmark code performs one forward and one backward pass through the HyperGaussian module, which outputs the conditional mean and the uncertainty.
3.3 FlashAvatar with HyperGaussians
@@
Our HyperGaussian representation can be integrated into existing 3DGS pipelines. To demonstrate this, we inject our conditional HyperGaussians into FlashAvatar [51]. FlashAvatar deploys a deformation MLP $F_{\theta}$, which maps FLAME [27] expression parameters $\boldsymbol{\psi}$ to per-Gaussian offsets $\Delta\boldsymbol{\mu}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{r}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{s}_{\boldsymbol{\psi}}$ for position, rotation, and scale. These offsets are then applied to the mesh-attached Gaussians to help model fine details. The deformation MLP further relies on auxiliary input consisting of positional encodings [34] of canonical mesh positions $\boldsymbol{\mu}_{T}$.
@@
Fig. 2 illustrates how we replace 3DGS with HyperGaussians. We modify the deformation MLP to output a per-Gaussian latent $\boldsymbol{z}_{\boldsymbol{\psi}}$ instead of offsets. We then compute the conditional distributions $p(\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}})$, $p(\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}})$ and $p(\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}})$ for every Gaussian. Recall from Sec. 3.2 that the conditional means and covariance matrices can be computed efficiently in closed form. The conditional means are then fed to the rest of the pipeline analogous to the offsets $\Delta\boldsymbol{\mu}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{r}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{s}_{\boldsymbol{\psi}}$ from FlashAvatar. More specifically, the parameters in global space are obtained via
$\begin{split}\boldsymbol{\mu}_{\boldsymbol{\psi}}&=\boldsymbol{\mu}_{M}\oplus\mathbb{E}\bigl{[}\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}}\bigr{]}\\ \boldsymbol{r}_{\boldsymbol{\psi}}&=\boldsymbol{r}\oplus\mathbb{E}\bigl{[}\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}}\bigr{]}\\ \boldsymbol{s}_{\boldsymbol{\psi}}&=\boldsymbol{s}\oplus\mathbb{E}\bigl{[}\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}}\bigr{]},\end{split}$ (7)
where $\boldsymbol{\mu}_{\boldsymbol{\psi}}$ are the positions on the posed mesh and $\boldsymbol{r}$, $\boldsymbol{s}$ are base rotation and scale. We find a latent dimensionality of $n=8$ to perform best for face avatars, but already a single latent dimension ($n=1$) improves over the baseline in Tab. 3.
Our method improves the final quality for thin structures like the glass frames, teeth, and specular reflection in Fig. 1 and Fig. 7. For further details about FlashAvatar, please refer to Xiang et al. [51].
Figure 6: Uncertainty Quantification (Sec. 4.2) on one of the training subjects. Green denotes low uncertainty, while Red denotes high uncertainty. Note that the semantic structure arises purely from the formulation without additional supervision.
@@ 4.1 Setting
Dataset
We compare face avatars trained on subjects from $4$ datasets. For quantitative results in Tab. 2 and qualitative comparisons in Figs. 7 and 8, we compare on 5 datasets from previous works [61, 58, 14, 12, 11], consisting of a total of 19 subjects. All videos are sub-sampled to 25 FPS and resized to a resolution of $512\times 512$. The length of the videos varies between $1$ and $3$ minutes, and we train on $2000$ frames and use the last $500$ frames for testing, following related works [51].
Preprocessing
@@
Our case study on FlashAvatar [51] builds directly on top of their public PyTorch [40] codebase. We inherit all losses and hyperparameters as shown in Sec. 3.3 and set the learning rate for our HyperGaussian parameters to $\eta=10^{-4}$. For each video, we train for $00030^{\prime}$ 000 iterations, which results in roughly $10$ epochs and takes between $15$ and $20$ minutes on a 24GB NVIDIA RTX 4090.
4.2 Comparisons
Fig. 7 shows a qualitative comparison for self-reenactment. SplattingAvatar [46] is not designed to represent expression-dependent appearance effects. This leads to unrealistic renderings for reflective regions like eyes and glasses, and artifacts in regions with strong deformations like the mouth and the eyes. Note how it fails to close the eyelid in the second row. MonoGaussianAvatar [8] predicts expression-dependent offsets but struggles with specular reflections. FlashAvatar uses an MLP to predict expression-dependent Gaussian parameter offsets to optimized 3D Gaussians. Their output lacks detail for thin structures like the glass frames or gaps between teeth, and it does not work well for specular reflections on the eyes and glasses. Ours handles such high-frequency details more gracefully. Keep in mind that the only difference between Ours and FlashAvatar is the Gaussian representation. Without any other changes to the architecture and no hyperparameter tuning, HyperGaussian is capable of rendering more accurate specular reflections and thin structures in the mouth, eyes, and glass frames.
We qualitatively compare cross-reenactment in Fig. 8. Note the distorted mouth and blurry hair in SplattingAvatar. MonoGaussianAvatar renders a more realistic face but exhibits unrealistic deformations on the jaw. FlashAvatar can render the correct expression, but the teeth and hair appear blurry. Ours produces higher-quality renders showing individual teeth and sharper hair. This can also be observed in a side-by-side comparison for training convergence in Fig. 3. Our HyperGaussian leads to a faster conversion than vanilla 3DGS.
Finally, it’s important to mention that MonoGaussianAvatar is substantially heavier than FlashAvatar and Ours. MonoGaussianAvatar requires over 12 hours of training time and uses over 100,000 Gaussians. FlashAvatar and Ours are much more lightweight with only $14,876$ Gaussians and less than 20 minutes training time. Despite the much longer training and the higher number of Gaussians for related works, our HyperGaussians outperforms them both. We encourage the reader to check out the supplementary HTML page to see animated results.
Method PSNR ↑ SSIM$\bigl{(}10^{-1}\bigr{)}$ LPIPS$\bigl{(}10^{-2}\bigr{)}$
@@ 4.2 Comparisons
FlashAvatar [51] 29.43 9.466 5.107
Ours 29.99 9.510 4.978
Table 2: Quantitative comparison with state-of-the-art digital avatar reconstruction methods from monocular video across 19 subjects from 5 datasets [61, 58, 14, 12, 11]. Ours corresponds to FlashAvatar with 8-dimensional HyperGaussians without any other modifications. Please see Sec. 3.3 for details.
Uncertainty from Covariances
We observe an emerging property about HyperGaussians, which arises naturally throughout training. HyperGaussians are at their core multivariate Gaussian distributions. Their conditional covariance matrices indicate the variance of each Gaussian across the different expressions of the training subject and can be intuitively interpreted as uncertainty. We compute the uncertainties $\sigma:=\log\det\boldsymbol{\Sigma}_{a|b}=-2\operatorname{tr}\log\boldsymbol{L}_{aa}$ for each conditional distribution $p(\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}})$, $p(\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}})$, and $p(\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}})$. Please refer to Eq. 14 in the Sec. A.1 for details. We visualize an example in Fig. 6. Red indicates high uncertainty, and green indicates low uncertainty. This agreement between the uncertainty estimates and what would intuitively be considered difficult regions is an inductive bias of our formulation and does not require explicit supervision.

### v3

@@ 3.2 HyperGaussians
HyperGaussians are an extension of 3D Gaussians to higher dimensions. As an example, a vanilla Gaussian in Kerbl et al. [26] has $m=3$ dimensions for the mean $\boldsymbol{\mu}_{\mathrm{3D}}$ representing its position. Intuitively, our HyperGaussian generalizes the vanilla Gaussian primitive to $(m+n)$ dimensions. We call $m$ the attribute dimensionality and the additional dimensions $n$ the latent dimensionality. In the following, we use subscripts $a$ and $b$ to indicate general random quantities. In practice, we identify them with concrete 3D Gaussian attributes and local latent embeddings, respectively. We later show that high-dimensional extensions of 3DGS admit a probabilistic interpretation.
@@
This paragraph formally describes HyperGaussians. Consider a general random vector $\boldsymbol{\gamma}\sim\mathcal{N}\bigl(\boldsymbol{\mu},\boldsymbol{\Sigma}\bigr)$} that is partitioned into vectors $\boldsymbol{\gamma}=(\boldsymbol{\gamma}_{a},\boldsymbol{\gamma}_{b})^{\top}$}, where $\boldsymbol{\gamma}_{a}\in\mathbb{R}^{m}$ and $\boldsymbol{\gamma}_{b}\in\mathbb{R}^{n}$.
@@
with $\boldsymbol{\Sigma}_{ba}=\boldsymbol{\Sigma}_{ab}^{\top}$. In our model, $\boldsymbol{\gamma}_{a}$ is instantiated with 3D Gaussian attributes $\mathcal{A}$ (e.g., position $\mathcal{A}_{\boldsymbol{\mu}_{\mathrm{3D}}}$) and $\boldsymbol{\gamma}_{b}$ with a latent code $\boldsymbol{z}$. Following NDGS [13], we parameterize the mean $\boldsymbol{\mu}$ directly and decompose the covariance $\boldsymbol{\Sigma}$ into its Cholesky factor $\boldsymbol{L}$, such that $\boldsymbol{\Sigma}=\boldsymbol{L}\boldsymbol{L}^{\top}$, where $\boldsymbol{L}$ is a lower triangular matrix with positive diagonal entries. Training requires optimizing the parameters of the Cholesky factor $\boldsymbol{L}$ together with the mean $\boldsymbol{\mu}$. Please read on to the end of this section for a detailed explanation of which parameters require training.
@@
HyperGaussians can be splatted with the differentiable rasterizer proposed by Kerbl et al. [26] after being reduced to the attribute dimensionality. This can be done by conditioning on the latent dimensions. Geometrically, conditioning corresponds to taking an $m$-dimensional slice through the multivariate Gaussian with $(m+n)$ total dimensions. More formally, we compute the conditional distribution $p(\boldsymbol{\gamma}_{a}|\boldsymbol{\gamma}_{b})=\mathcal{N}\bigl(\boldsymbol{\mu}_{a|b},\boldsymbol{\Sigma}_{a|b}\bigr)$}. Since HyperGaussians follow a multivariate Gaussian distribution, the conditional distribution can be computed in closed form [2]:
@@
We use the mean of this conditional distribution to recover a vanilla 3D Gaussian, which can be splatted with the differentiable rasterizer introduced by Kerbl et al. [26]. Note that for HyperGaussians, the conditional covariance matrix is unused during splatting, contrary to NDGS [13], and instead offers a way to derive probabilistic quantities, such as uncertainty (see the supp. mat.).
Bayesian Interpretation
A key differentiation between HyperGaussians and existing multi-dimensional Gaussians [13, 18, 19] is that our formulation is strictly more expressive. To show this, we interpret high-dimensional variants of 3DGS as probabilistic descriptors of their underlying 3D primitives. More formally, they model 3D Gaussian attributes $\mathcal{A}$ as random variables that are coupled to a random latent vector $\boldsymbol{z}$ via a joint multivariate normal distribution $p(\mathcal{A},\boldsymbol{z})$. This relationship is learned during training. Due to the properties of multivariate Gaussians, the coupling induces a posterior density $p(\mathcal{A}|\boldsymbol{z})$ which is again Gaussian. Computing the conditional mean (i.e., the reduction to 3D primitives for splatting) is akin to maximum a posteriori (MAP) estimation, since
$\mathbb{E}[\mathcal{A}|\boldsymbol{z}]=\operatorname*{\arg\max}_{\mathcal{A}}p(\mathcal{A}|\boldsymbol{z})$ (4)
for Gaussian distributions. Most importantly, the associated Gaussian prior $p(\mathcal{A})$ acts as an implicit regularizer, which assists in preserving details during test-time inference.
We therefore identify NDGS-based methods as well-behaved distributions over 3DGS scenes. To the best of our knowledge, we are the first to highlight this connection and provide theoretical justification for high-dimensional Gaussian Splatting. This view further allows us to analyze current limitations of NDGS [13], which HyperGaussians aim to mitigate. In particular, NDGS uses $\boldsymbol{\mu}_{\mathrm{3D}}\leftarrow\mathbb{E}[\mathcal{A}_{\boldsymbol{\mu}_{\mathrm{3D}}}|\boldsymbol{z}]$}, $\boldsymbol{\Sigma}_{\mathrm{3D}}\leftarrow\mathrm{Cov}[\mathcal{A}_{\boldsymbol{\mu}_{\mathrm{3D}}}|\boldsymbol{z}]$}, and $p(\mathcal{A}_{\boldsymbol{\mu}_{\mathrm{3D}}},\boldsymbol{z})$ to modulate the opacity. In other words, NGDS only models the dependence between the position and the latent code, while the scaling and rotation are independent because the conditional covariance matrix does not consider the realization of $\boldsymbol{z}$ (see Eq. 3, $\boldsymbol{z}$ corresponds to $\gamma_{b}$). On the other hand, HyperGaussians additionally model $p(\mathcal{A}_{\boldsymbol{s}}|\boldsymbol{z})$ and $p(\mathcal{A}_{\boldsymbol{r}}|\boldsymbol{z})$, allowing the 3D primitives to scale and rotate according to the latent code. This means that our parameterization accounts for conditional distributions of 3DGS scenes with strictly larger support than NDGS.
@@
A naïve implementation of the conditioning in Eq. 3, as in NDGS [13], is very inefficient for large latent codes $\boldsymbol{\gamma}_{b}$. The bottleneck lies in storing and inverting the conditional covariance matrix $\boldsymbol{\Sigma}_{bb}\in\mathbb{R}^{n\times n}$. The latent dimensionality $n$ is typically much larger than the dimensionality after conditioning $m$ ($n\gg m$).
@@
We propose the inverse covariance trick that substantially reduces the cost of conditioning from $\mathcal{O}\bigl(n^{3}+mn^{2}\bigr)$ to $\mathcal{O}\bigl(m^{3}+m^{2}n\bigr)$, a reduction from cubic to linear growth with respect to the dominating factor nn! The key idea is to reformulate the HyperGaussians in terms of their precision matrix $\boldsymbol{\Lambda}=\boldsymbol{\Sigma}^{-1}$ such that $\boldsymbol{\gamma}\sim\mathcal{N}\bigl(\boldsymbol{\mu},\boldsymbol{\Lambda}^{-1}\bigr)$. Consider the general block matrix view as in Eq. 2:
$\boldsymbol{\Sigma}^{-1}=\boldsymbol{\Lambda}=\begin{bmatrix}\boldsymbol{\Lambda}_{aa}&\boldsymbol{\Lambda}_{ab}\\ \boldsymbol{\Lambda}_{ba}&\boldsymbol{\Lambda}_{bb}\end{bmatrix}$, (5)
with $\boldsymbol{\Lambda}_{ba}=\boldsymbol{\Lambda}_{ab}^{\top}$. The conditional mean can now be expressed as
$\boldsymbol{\mu}_{a|b}=\boldsymbol{\mu}_{a}-\boldsymbol{\Lambda}_{aa}^{-1}\boldsymbol{\Lambda}_{ab}\bigl(\boldsymbol{\gamma}_{b}-\boldsymbol{\mu}_{b}\bigr).$ (6)
And the covariance matrix simply as $\boldsymbol{\Sigma}_{a|b}=\boldsymbol{\Lambda}_{aa}^{-1}$. This new formulation only requires storing and inverting the much smaller block $\boldsymbol{\Lambda}_{aa}$ of the precision matrix. While this change seems minor in the derivation, it significantly improves both speed and memory usage. Note that it is not necessary to optimize the full precision matrix $\boldsymbol{\Lambda}$, it suffices only to parameterize $\boldsymbol{\Lambda}_{aa}$ and $\boldsymbol{\Lambda}_{ab}$. This significantly reduces the #\#parameters from $\mathcal{O}\bigl((m+n)^{2}\bigr)$ to $\mathcal{O}\bigl(m^{2}+mn\bigr)$. See Fig. 3 for a benchmark comparison and the supp. mat. for more details, where we also describe how to further reduce the runtime to $\mathcal{O}\bigl(m^{2}n\bigr)$ by exploiting the readily available Cholesky decomposition of $\boldsymbol{\Lambda}_{aa}$.
@@
Figure 3: Benchmark Results on conditioning for ${\sim}15$k HyperGaussians with attribute dimension $m=3$ (e.g., position) and varying latent dimension $n$. We average the measurements across 1000 runs on an NVIDIA GeForce RTX 2080 Ti after an initial warm-up. The benchmark code performs one forward and one backward pass.
3.3 Case Studies
HyperGaussians can be integrated into existing 3DGS pipelines and improve their quality. To demonstrate this, we inject our conditional HyperGaussians into FlashAvatar [70] and GaussianHeadAvatar [71]. In the following, we briefly describe these two case studies. Please see the supp. mat. for additional details.
Integrating HyperGaussians into FlashAvatar
@@ 3.3 Case Studies
FlashAvatar deploys a deformation MLP $F_{\theta}$, which maps FLAME [38] expression parameters $\boldsymbol{\psi}$ to per-Gaussian offsets $\Delta\boldsymbol{\mu}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{r}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{s}_{\boldsymbol{\psi}}$ for position, rotation, and scale. Note that we adopt the notation from FlashAvatar in this section to ensure consistency. These offsets are then applied to the mesh-attached Gaussians to help model fine details. The deformation MLP further relies on auxiliary input consisting of positional encodings [47] of canonical mesh positions $\boldsymbol{\mu}_{T}$.
@@
Fig. 2 illustrates how we replace 3DGS with HyperGaussians. We modify the deformation MLP to output a per-Gaussian latent $\boldsymbol{z}_{\boldsymbol{\psi}}$ instead of offsets. We then compute the conditional distributions $p(\Delta\boldsymbol{\mu}|\boldsymbol{z}_{\boldsymbol{\psi}})$, $p(\Delta\boldsymbol{r}|\boldsymbol{z}_{\boldsymbol{\psi}})$ and $p(\Delta\boldsymbol{s}|\boldsymbol{z}_{\boldsymbol{\psi}})$ for every Gaussian. Recall from Sec. 3.2 that the conditional means and covariance matrices can be computed efficiently in closed form. The conditional means are then fed to the rest of the pipeline, exactly like the offsets $\Delta\boldsymbol{\mu}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{r}_{\boldsymbol{\psi}}$, $\Delta\boldsymbol{s}_{\boldsymbol{\psi}}$ from FlashAvatar.
We find that a latent dimensionality of $n=8$ performs best for face avatars, but even a single latent dimension ($n=1$) improves over the baseline (see the supp. mat. for an ablation). Figs. 1 and 4 show the effect of HyperGaussians on thin structures like the glass frames, teeth, and specular reflection.
Integrating HyperGaussians into GaussianHeadAvatar
To further show the capabilities of HyperGaussians for enhancing existing 3DGS pipelines, we integrate them into the state-of-the-art avatar method for multi-view videos, GaussianHeadAvatar [71]. GaussianHeadAvatar models the head using expression-conditioned 3D Gaussians with multi-channel colors. More specifically, a dynamic generator takes expression and pose coefficients of a 3DMM and predicts per-Gaussian offsets with respect to a canonical point cloud. The rendered feature image is then passed through a super-resolution module to obtain the final render in RGB space. We modify their architecture to predict latents for our HyperGaussians rather than offsets. Similar to our FlashAvatar modification, we condition the HyperGaussians to obtain regular 3D Gaussians that can be splatted.
Our new representation greatly improves deformation effects, such as wrinkles. Also, it produces sharper reconstructions in general, but particularly in the eyes and glass frames (Figs. 1 and 6). We kindly refer to Xu et al. [71] and the supp. mat. for details on the baseline method.
@@ 4.1 Setting
Monocular Dataset
For quantitative results in Tab. 1 (top part) and qualitative comparisons in Figs. 4 and 5, we compare on 5 datasets from previous works [79, 76, 21, 17, 15], consisting of a total of 19 subjects. All videos are subsampled to 25 FPS and resized to $512\times 512$. The videos range from $1$ to $3$ min in length; we train on $2000$ frames and use the last $500$ for testing, following related work [70].
@@
Multi-view Dataset
We evaluate the effectiveness of our method applied to GaussianHeadAvatar [71] on $10$ subjects from the NeRSemble [28] dataset, shown in Tab. 1 (bottom part) and Fig. 6. Following GaussianHeadAvatar, the videos are sampled at 25 FPS, resized, and cropped to $2048\times 2048$. We hold out the FREE sequence for all subjects for testing and train on all remaining sequences. This results in ${\sim}2500$ training frames and ${\sim}300$ test frames, each viewed from $16$ cameras.
For preprocessing, we follow GaussianHeadAvatar and use off-the-shelf tools for background removal [40] and 2D landmark detection [8]. We then track a BFM [20] model using the landmark annotations.
@@
Our case studies on FlashAvatar [70] and GaussianHeadAvatar [71] build directly on top of their public PyTorch [54] codebases. We inherit all losses and hyperparameters as shown in Sec. 3.3 and set the learning rate for our HyperGaussian parameters to $\eta=10^{-4}$. For FlashAvatar, we train on each video for 3030k iterations, which takes between $15$ and $20$ min on an NVIDIA RTX 4090. For GaussianHeadAvatar, we train on each subject for 600600k iterations, which takes over $2$ days. We publicly release an implementation of HyperGaussians to streamline integration into other codebases.
[…]
@@
FlashAvatar (FA) [70] 29.43 9.466 5.107
Ours (FA) 29.99 9.510 4.978
GaussianHeadAvatar (GHA) [71] 24.10 8.819 20.273
Ours (GHA) 24.38 8.819 19.768
Table 1: Quantitative comparison with state-of-the-art digital avatar reconstruction methods from monocular video across 19 subjects from 5 datasets [79, 76, 21, 17, 15] (top) and from multi-view video across 10 subjects from NeRSemble [28] (bottom). Ours (FA) and Ours (GHA) correspond to FlashAvatar and GaussianHeadAvatar with 8-dimensional HyperGaussians without any other modifications. Please see Secs. 3.3 and 3.3 for details.
4.2 Comparisons
Monocular Setting
Fig. 4 shows a qualitative comparison for self-reenactment. SplattingAvatar [60] is not designed to represent expression-dependent appearance effects. This leads to unrealistic renderings in reflective regions, such as eyes and glasses, and to artifacts in areas with strong deformations, such as the mouth and eyes. Note how it fails to close the eyelid in the second row. MonoGaussianAvatar [10] predicts expression-dependent offsets but struggles with specular reflections. FlashAvatar uses an MLP to predict expression-dependent Gaussian parameter offsets. Their output lacks detail for thin structures, such as the glass frames or gaps between teeth, and it does not handle specular reflections on the eyes and glasses well. Ours handles such high-frequency details more gracefully. It is crucial to keep in mind that the only difference between Ours and FlashAvatar is the Gaussian representation. Without any other changes or hyperparameter tuning, HyperGaussians are capable of rendering more accurate specular reflections and thin structures in the mouth, eyes, and glass frames.
We qualitatively compare cross-reenactment in Fig. 5. Note the distorted mouth and blurry hair in SplattingAvatar. MonoGaussianAvatar renders a more realistic face but exhibits unrealistic jaw deformations. FlashAvatar can render the correct expression, but the teeth and hair appear blurry. Ours produces higher-quality renders showing individual teeth and sharper hair. This can also be observed in a side-by-side comparison for training convergence (see supp. mat.).
Finally, it is important to mention that MonoGaussianAvatar is substantially heavier than FlashAvatar and Ours. MonoGaussianAvatar requires over 1212h of training time and uses over 100100k Gaussians. FlashAvatar and Ours are much lighter, with only ${\sim}15$k Gaussians and less than 2020min of training time. Despite this, our HyperGaussians outperforms them both. We encourage readers to visit the supplementary HTML page to see animated results.
Multi-view Setting
We further evaluate HyperGaussians on multi-view avatar reconstruction. By replacing 3D Gaussians with our HyperGaussians, we significantly improve the visual quality of GaussianHeadAvatar (Fig. 6). Our more expressive representation enables the model to faithfully reconstruct challenging nonlinear effects. This is evident in rows 1, 3, and 4, where our method improves the appearance of specular highlights on eyes, teeth, and glasses. Moreover, the rich local context of HyperGaussians allows for accurate modeling of complex skin deformations, which the baseline failed to capture entirely (rows 2 and 4).
Note that, thanks to the inverse covariance trick, the HyperGaussians integration results in very little overhead. Specifically, the training time increased by only 3030min ($1\%$), which is negligible considering the total of over 22d.
Figure 6: Enhancing GaussianHeadAvatars [71] with HyperGaussians boosts high-frequency details like wrinkles and reflections in the eyes, glasses, and teeth. See Tab. 1 for metrics.